OpenStudy (anonymous):

Find the interest rate that is being compounded every three months which is equivalent to an effective annual rate of 4%.

OpenStudy (solomonzelman):

\[A=P(1+\frac{r}{n})^{nt}\]A is the result after the investment P is the initial investment r is the interest rate n is number of times that it's compound per year t is the number of years

OpenStudy (wolf1728):

Compounded Rate = [(1 + r/n)^n]-1 'r' is rate before compounding 'n' - number of compounding periods Compounded Rate = (1+.04/4)^4)-1 Compounded Rate = (1.01)^4 -1 Compounded Rate = 1.04060401 -1 Compounded Rate = .04060401 Compounded Rate = 4.060401 Sources: Formula http://1728.org/rate.htm Calculator http://1728.org/intrate.htm